# Transformation of orientation and rotation angles of synchronous satellites: Application to the Galilean moons

### v1

@@ 5 Dynamical models
The Galilean satellites are assumed to be locked in the Cassini state, with their rotation variations described as a combination of variations in the rotation angle (librations, as defined in Section 4.3), variations in the orientation of the spin axis in space (nutations, as defined in Section 3.2), and variations in the orientation of the spin axis relative to the Body Frame (polar motion). These rotation variations are caused by the external gravitational torque exerted by Jupiter on the flattened shape of the satellites. In the case of Titan, in addition to the external torque exerted by Saturn, angular momentum exchanges between the atmosphere and lakes and the solid surface also contribute to the rotation variations (the resulting variations in rotation around the spin axis are then called length-of-day variations). In this section, we identify relevant dynamical models to create Euler angle solutions to be transformed into equatorial coordinates in Section 6.3. Since we provide a transformation method for series of rotation variations, we are only interested here in rotation models whose solution is expressed in series.
Series for rotational variations can be modeled within the angular momentum (AM) framework, by solving equations that relate the rate of change of the rotation AM to the external torques applied. The AM equations can be solved numerically, before a frequency analysis is applied to the solution (e.g. Rambaux et al. [2011]). Fitted series for the rotational variations can also be obtained within an Hamiltonian framework (e.g. Noyelles [2009]). Another way of approaching the problem is to solve the AM equations analytically, using a semi-analytical orbital theory to express the forcing, so that the solution is directly expressed in series (e.g. Bills 2005, Baland et al. 2012, Yseboodt and Van Hoolst 2014). The latter approach requires a number of working assumptions, particularly with regard to the quantities and terms that can be neglected or not. Such a series obtained from an analytical solution will generally contain fewer terms than a series fitted from a numerical solution. Nevertheless, by extending an analytical solution to a sufficient order, it should be possible to recover a sufficient number of terms and tend towards the numerical solution. For example, Coyette et al. [2025] recently extended an analytical AM theory to the second order in small quantities and recovered rotational variation terms that had previously only been identified from a frequency analysis. Here we will provide simplified dynamical models (analytical AM models developed to the first order in small quantities) identified in the literature. Providing a very detailed theory of the rotation of Galilean satellites is outside the scope of the present study and is not necessary to assess the accuracy of the transformations defined in Sections 3 and 4.
@@ 5.1 Spin orientation model
an expression that is consistent with the majority of the literature (e.g. Noyelles 2010, Henrard and Schwanen 2004, Boué 2020, Coyette et al. 2025). Bills [2005] and Bills and Nimmo [2011] obtain different solutions, due to their incorrect definitions of the coupling constant $\kappa$. The exact expression of Eq. (39) has been demonstrated in Appendix A of Baland et al. [2012].
@@
In the limit of small orbital obliquity and neglecting tidal dissipation, Eqs. (42) correct Eqs. (E4-E5) of Bills and Scott [2022] as obtained from their Eq. 4.8 (with $\boldsymbol{\hat{\textbf{s}}}.\boldsymbol{\hat{\textbf{n}}}\sim 1$ and $p_{3}=0$). Eqs. (42) can be integrated numerically to provide a benchmark for assessing the accuracy of the transformation presented in section 3.
For the case where the satellites harbor an internal global liquid layer, as will be considered in Section 8.2, we will use the solution presented in Appendix 5 of Baland et al. [2019], which extends the model of Baland et al. [2011, 2012] by including the hydrodynamic pressure (Poincare flow) at the interfaces between the ocean and the solid layers. That model assumes that the solid layers behave rigidly.
@@ 5.2 Libration model
where $\boldsymbol{\textbf{$\Omega$}}=(\omega_{x},\omega_{y},\omega_{z})$ is the rotation vector, $\boldsymbol{\textbf{N}}=(N_{x},N_{y},N_{z})$ is the total torque exerted on the satellite, and L is the angular momentum. For a fully rigid triaxial satellite, $L_{z}=C\,\omega_{z}$ with $\omega_{z}\simeq\dot{\psi}+\dot{\phi}_{Euler}=\dot{\phi}_{Inertial}$ at first order in $\theta$ and the libration equation reads, see also Eqs. (5.48) of Murray and Dermott [2000]:
$C\,\dot{\omega_{z}}-(A-B)\,\omega_{x}\,\omega_{y}=N_{z}.$ (44)
In a simple first order theory, the product of the polar motion components $\omega_{x}$ and $\omega_{y}$ can be neglected and $N_{z}$ approximated at first order in small quantities, so that
$C\,\ddot{\phi}_{Inertial}=-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin 2\left(\phi_{Inertial}-l_{JupEq}\right),$ (45)
@@
$C\,\ddot{\phi}_{pericenter}=-\frac{3}{2}n^{2}(B-A)\,\left(\frac{a}{d}\right)^{3}\sin 2\left(\phi_{pericenter}-f\right)\qquad\textrm{[For a Keplerian orbit]},$ (46)
with $f$ the true anomaly, see e.g. Eq. (1) of Van Hoolst et al. [2008]. The right-hand side of the equation can first be expressed as series in eccentricity $e$ before solving the dynamical equation. The solution consists in a main diurnal librations and smaller terms at harmonics of the orbital period.
Orbital perturbations lead to changes in the gravitational torque of Jupiter on the satellite compared to the torque in the Keplerian case. Based on the chosen orbital theory, the factor of $\sin 2(\phi_{Inertial}-l_{JupEq})/d^{3}$ can first be expressed as series before solving the dynamical equation. Depending on the chosen definition for the libration angle (with respect to the forcing $\gamma_{f}$ or with respect to the uniform rotation $\gamma_{u}$), Eq. (45) can take two different forms:
$C\,(\ddot{\gamma}_{f}+\ddot{\mathcal{L}})$ $=$ $-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin(2\,\gamma_{f}+2\,M-2\,f),$ (47a)
$C\,\ddot{\gamma}_{u}$ $=$ $-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin(2\,\gamma_{u}+2\,L-2\,l_{JupEq}).$ (47b)
@@ 6.1 Spin orientation series for solid Galilean satellites, in $(s_{x},s_{y})$
Following the decomposition described in Section 2.1, the orbit normal coordinates ($n_{x},n_{y})$ are written as the sum of a linear term (with the $tr_{x}$ and $tr_{y}$ rates, second order terms by construction) and of a trigonometric series (noted $\Delta$, and of first order by construction):
@@
Table 2: Values for the mass ($M$), radius ($R$), mean motion ($n_{L}$), normalized principal moments of inertia $A/MR^{2}$ and $C/MR^{2}$, and free precession period ($2\pi/\omega_{f}$) of the four Galilean satellites. The moments of inertia AA and $C$ are derived from the observed $J_{2}$ and $C_{22}$ gravitational coefficients and from the mean moment of inertia II (see Anderson et al. [2001] for Io, Gomez Casajus et al. [2021] for Europa and Schubert et al. [2004] for Ganymede and Callisto). The $n_{L}$ values are computed from JUP387 and differ from those in Lainey et al. [2006] by less than $3\,10^{-8}$ rad/d. The last column evaluates the amplitude at the satellite surface of an arbitrary angular displacement of $1^{\circ}$.
@@
Io $2.8\,10^{-5}$ $2.4\,10^{-5}$ $6.5\,10^{-5}$ $5.9\,10^{-5}$ $203.488958424$ $1.4\;10^{-5}$
Europa $2.8\,10^{-5}$ $2.4\,10^{-5}$ $6.4\,10^{-5}$ $5.9\,10^{-5}$ $101.374724491$ $2.49\,10^{-3}$
Ganymede $2.8\,10^{-5}$ $2.5\,10^{-5}$ $6.5\,10^{-5}$ $5.9\,10^{-5}$ $50.317607524$ $4.7\;10^{-4}$
Callisto $2.8\,10^{-5}$ $2.7\,10^{-5}$ $6.5\,10^{-5}$ $5.9\,10^{-5}$ $21.571072373$ $1.47\,10^{-3}$
Table 3: Trends in the various quantities considered. The $tr_{x}$ and $tr_{y}$ trends, obtained from the JUP387 orbital theory, apply to both the orbit normal and spin axis coordinates $n/s_{x}$ and $n/s_{y}$, respectively. The $tr_{\alpha}$, $tr_{\delta}$ and $tr_{\mu}$ trends are obtained after the transformation described in Eqs. (10) and (23b). The rotation rate in $W$ is obtained from Eq. (32). The last column shows the small increment in $\delta_{S}$ due to the power reduction of a squared sine term in the second-order part of Eq. (8b).
Figure 2: Temporal evolution of the spin axis coordinates in the Laplace plane $s_{x}$ and $s_{y}$ for the four Galilean satellites for a solid satellite. The time coverage is different for each satellite. The purple point is the Laplace pole, for which by definition $s_{x}=s_{y}=0$. The black dots in the right graphs are the J2000 spin positions.
@@
In Table 4, we have considered a truncation threshold of the order of $0.0001^{\circ}$ on the $i_{j}$, corresponding to the order of magnitude of the residuals in the orbit orientation angles between JUP387 and its frequency decomposition. We thus obtain $6$, $10$, 99, and $8$ terms for Io, Europa, Ganymede, and Callisto, respectively. The arguments (frequencies and phases) of the series result from the orbital dynamics and are combinations of the fundamental arguments associated with the Jovian system (see last columns). However, arguments identified as being the same combination may have a slightly different value for each satellite, due to the fit to a numerical integration. The precision on the $i_{j}$, $\varepsilon_{j}$ and $\theta_{j}$ as given in Table 4 is $0.0001^{\circ}$ ($0.00001$ rad/y and $0.001^{\circ}$ in frequency and phase, respectively). The typical precision of the time-series for $s_{x/y}(t)$ reconstructed from Table 4 over the 190021001900-2100 interval is approximately $0.0003^{\circ}$ (see Table 5). This corresponds to a relative truncation error below $1\%$. Neglecting terms related to trends would lead to typical errors of about $0.003^{\circ}$ after $100$ years in $s_{x/y}(t)$, greater than the truncation error.
@@
Table 4: $(\Delta n_{x},\Delta n_{y})$ and $(\Delta s_{x},\Delta s_{y})$ series, according to the notations of Eqs. (48-50). Inclination, orbital and inertial obliquity amplitudes $i_{j}$, $\varepsilon_{j}$ and $\theta_{j}$ at the frequency and phase $f_{j}$ and $\varphi_{j}$ obtained from the JUP387 orbital theory and assuming that the satellites are entirely solid and rigid. These series can be used directly to reconstruct time series for $s_{x/y}(t)$, and indirectly to reconstruct the time series for the angles $\theta(t)$, $\varepsilon(t)$, $\psi(t)$, and $\zeta(t)$ (see Section 6.2). The phases $\varphi_{j}$ are given with respect to J2000 epoch. The last columns show the identification to the fundamental arguments of Lainey et al. [2006] and the fundamental IAU WG frequencies $\dot{J}_{i}$.
@@
For each satellite, the orbital precession is dominated by the term with $j=1$ whose frequency corresponds to the node longitude proper frequency of that satellite. This may seem obvious, but note that this is not the case for all of the large satellites of Uranus, which strongly perturb each other [Baland et al., 2025]. The ratio $i_{2}/i_{1}$ is of $28\%$, $5\%$, $21\%$, and $12\%$, for Io, Europa, Ganymede, and Callisto by respectively. The dominant terms of spin precession are not necessarily the same as the dominant terms of orbital precession. The amplitude $\varepsilon_{j}$ can reach a value such that $\theta_{j}$ is large or very small, depending on the ratio of $\omega_{f}/\dot{\Omega}_{j}$ with $\omega_{f}=\kappa/C$ the natural frequency of the free retrograde precession. The free mode periods are given in Table 2. In the limit case $\omega_{f}/\dot{\Omega}_{j}\rightarrow 0$ (relatively short forcing period), $\varepsilon_{j}$ tends towards $i_{j}$ and $\theta_{j}$ tends towards 0 (meaning that the spin axis tends to align with the Laplace pole, e.g. term $j=3$ for Callisto). For $\omega_{f}/\dot{\Omega}_{j}\rightarrow\pm\infty$ (relatively long forcing period), $\varepsilon_{j}$ tends towards 0 and $\theta_{j}$ tends towards $i_{j}$ (meaning that the spin axis tends to follow the precessing orbit normal, e.g. term $j=5$ for Io). When the ratio $\omega_{f}/\dot{\Omega}_{j}$ is close to 1-1, $\varepsilon_{j}$ and $\theta_{j}$ can be resonantly amplified (e.g. term $j=3$ for Ganymede). The ratio $\theta_{2}/\theta_{1}$ is of $26\%$, $5\%$, $18\%$, $16\%$ for Io, Europa, Ganymede and Callisto, respectively. Note that $\theta_{3}/\theta_{1}=22\%$ for Ganymede, while $i_{3}/i_{1}$ is only of $9\%$. The largest orbital obliquity amplitudes is $\varepsilon_{1}$ for Io, Europa and Callisto, but $\varepsilon_{3}$ for Ganymede. The second largest orbital obliquity amplitude is $\varepsilon_{4}$, $\varepsilon_{5}$, $\varepsilon_{1}$, and $\varepsilon_{2}$ for Io, Europa, Ganymede, and Callisto, respectively.
@@
Table 5: Comparison of different errors affecting the angles over the 1900-2100 period. The first source of error arises from the truncation applied in Table 4. Each term in the $s_{x/y}$ series is truncated at a level of $0.0001^{\circ}$ in amplitude, which is also the truncation threshold applied to the series. The cumulative errors on $s_{x/y}(t)$, and therefore on $\theta(t)$, $\varepsilon(t)$, and $\zeta(t)$, lead to a total truncation error of about $0.0003^{\circ}$. The cumulative truncation errors on the node longitude $\psi$ reach up to $0.6^{\circ}$ for Io. The second source of error is the development error, arising when $\theta$, $\psi$, $\varepsilon$, and $\zeta$ are developed as series themselves. A third possible source of error stems from the omission of the linear and Poisson terms.

### v2

@@ 5 Dynamical models
The Galilean satellites are assumed to be locked in the Cassini state, with their rotation variations described as a combination of variations in the rotation angle (librations, as defined in Section 4.3), variations in the orientation of the spin axis in space (nutations, as defined in Section 3.2), and variations in the orientation of the spin axis relative to the Body Frame (polar motion). These rotation variations are caused by the external gravitational torque exerted by Jupiter on the flattened shape of the satellites. In this section, we identify the relevant dynamical models used to construct Euler angle solutions, which will then be transformed into equatorial coordinates in Section 6.3. Since we provide a transformation method for series of rotation variations, we are only interested here in rotation models whose solution is expressed in series. Note that although the models described here apply essentially to all synchronous natural satellites, in the case of Titan, in addition to the external torque exerted by Saturn, angular momentum exchanges between the atmosphere and lakes and the solid surface also contribute to the rotation variations (the resulting variations in rotation around the spin axis are then called length-of-day variations).
Series for rotational variations can be modeled within the angular momentum (AM) framework, by solving equations that relate the rate of change of the rotation AM to the external torques applied. The AM equations can be solved numerically, before a frequency analysis is applied to the solution (e.g. Rambaux et al. 2011). Fitted series for the rotational variations can also be obtained within an Hamiltonian framework (e.g. Noyelles 2009). Another way of approaching the problem is to solve the AM equations analytically, using a semi-analytical orbital theory to express the forcing, so that the solution is directly expressed in series (e.g. Bills 2005, Baland et al. 2012, Yseboodt and Van Hoolst 2014). The latter approach requires a number of working assumptions, particularly with regard to the quantities and terms that can be neglected or not. Such a series obtained from an analytical solution will generally contain fewer terms than a series fitted from a numerical solution. Nevertheless, by extending an analytical solution to a sufficient order, it should be possible to recover a sufficient number of terms and tend towards the numerical solution. For example, Coyette et al. [2026] recently extended an analytical AM theory to the second order in small quantities and recovered rotational variation terms that had previously only been identified from a frequency analysis. Here we will provide simplified dynamical models (analytical AM models developed to the first order in small quantities) identified in the literature. Providing a very detailed theory of the rotation of Galilean satellites is outside the scope of the present study and is not necessary to assess the accuracy of the transformations defined in Sections 3 and 4.
@@ 5.1 Spin orientation model
an expression that is consistent with the majority of the literature (e.g. Noyelles 2010, Henrard and Schwanen 2004, Boué 2020, Coyette et al. 2026). Bills [2005] and Bills and Nimmo [2011] obtain different solutions, due to their incorrect definitions of the coupling constant $\kappa$. The exact expression of Eq. (40) has been demonstrated in Appendix A of Baland et al. [2012].
@@
In the limit of small orbital obliquity and neglecting tidal dissipation, Eqs. (43) correct Eqs. (E4-E5) of Bills and Scott [2022] as obtained from their Eq. (4.8) (with $\boldsymbol{\hat{\textbf{s}}}.\boldsymbol{\hat{\textbf{n}}}\sim 1$ and $p_{3}=0$). Eqs. (43) can be integrated numerically to provide a benchmark for assessing the accuracy of the transformation presented in section 3.
For the case where the satellites harbor an internal global liquid layer, as will be considered in Section 8.2, we will use the solution presented in Appendix 5 of Baland et al. [2019], which extends the model of Baland et al. [2011, 2012] by including the hydrodynamic pressure (Poincaré flow) at the interfaces between the ocean and the solid layers. That model assumes that the solid layers behave rigidly.
@@ 5.2 Libration model
where $\boldsymbol{\textbf{$\Omega$}}=(\omega_{x},\omega_{y},\omega_{z})$ is the rotation vector, $\boldsymbol{\textbf{N}}=(N_{x},N_{y},N_{z})$ is the total torque exerted on the satellite, and L is the angular momentum. For a fully rigid triaxial satellite, $L_{z}=C\,\omega_{z}$ with $\omega_{z}\simeq\psi^{\prime}+\phi^{\prime}_{Euler}=\phi^{\prime}_{Inertial}$ at first order in $\theta$ and the libration equation reads (see also Eqs. (5.48) of Murray and Dermott 2000):
$C\,\omega^{\prime}_{z}-(A-B)\,\omega_{x}\,\omega_{y}=N_{z}.$ (45)
The notation is the time derivative. In a simple first order theory, the product of the polar motion components $\omega_{x}$ and $\omega_{y}$ can be neglected and $N_{z}$ approximated at first order in small quantities, so that
$C\,\phi^{\prime\prime}_{Inertial}=-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin 2\left(\phi_{Inertial}-l_{JupEq}\right),$ (46)
@@
$C\,\phi^{\prime\prime}_{pericenter}=-\frac{3}{2}n^{2}(B-A)\,\left(\frac{a}{d}\right)^{3}\sin 2\left(\phi_{pericenter}-f\right)\qquad\textrm{[For a Keplerian orbit]},$ (47)
with $f$ the true anomaly, see e.g. Eq. (1) of Van Hoolst et al. [2008]. The right-hand side of the equation can first be expressed as series in eccentricity $e$ before solving the dynamical equation. The solution consists of a main diurnal libration and smaller terms at harmonics of the orbital period.
Orbital perturbations lead to changes in the gravitational torque of Jupiter on the satellite compared to the torque in the Keplerian case, see for example Yseboodt and Van Hoolst [2014]. Based on any orbital theory, the factor of $\sin 2(\phi_{Inertial}-l_{JupEq})/d^{3}$ can first be expressed as series before solving the dynamical equation. Depending on the chosen definition for the libration angle (with respect to the forcing $\gamma_{f}$ or with respect to the uniform rotation $\gamma_{u}$), Eq. (46) can take two different forms:
$C\,(\gamma^{\prime\prime}_{f}+\mathcal{L}^{\prime\prime})$ $=$ $-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin(2\,\gamma_{f}+2\,M-2\,f),$ (48a)
$C\,\gamma^{\prime\prime}_{u}$ $=$ $-\frac{3}{2}(B-A)\,\frac{GM_{J}}{d^{3}}\sin(2\,\gamma_{u}+2\,L-2\,l_{JupEq}).$ (48b)
@@ 6.1 Spin orientation series for solid Galilean satellites, in $(s_{x},s_{y})$
Following the decomposition described in Section 2.1, the orbit normal coordinates ($n_{x},n_{y}$) are written as the sum of a linear term (with the $tr_{x}$ and $tr_{y}$ rates, second order terms by construction) and of a trigonometric series (noted $\Delta$, and of first order by construction, see Section 3.2):
@@
Table 2: Values for the mass ($M$), radius ($R$), mean motion ($n_{L}$), normalized principal moments of inertia $A/MR^{2}$ and $C/MR^{2}$, and free precession period ($2\pi/\omega_{f}$) of the four Galilean satellites. The moments of inertia AA and $C$ are derived from the observed $J_{2}$ and $C_{22}$ gravitational coefficients and from the mean moment of inertia II (see Anderson et al. [2001] for Io, Gomez Casajus et al. [2021] for Europa and Schubert et al. [2004] for Ganymede and Callisto). The $n_{L}$ values are computed from JUP387 and differ from those in Lainey et al. [2006] by less than $3\times 10^{-8}$ rad/d. The last column evaluates the amplitude at the satellite surface of an arbitrary angular displacement of $1^{\circ}$.
@@
Io $2.8\times 10^{-5}$ $2.4\times 10^{-5}$ $6.5\times 10^{-5}$ $5.9\times 10^{-5}$ $203.488958424$ $1.4\times 10^{-5}$
Europa $2.8\times 10^{-5}$ $2.4\times 10^{-5}$ $6.4\times 10^{-5}$ $5.9\times 10^{-5}$ $101.374724491$ $2.49\times 10^{-3}$
Ganymede $2.8\times 10^{-5}$ $2.5\times 10^{-5}$ $6.5\times 10^{-5}$ $5.9\times 10^{-5}$ $50.317607524$ $4.7\times 10^{-4}$
Callisto $2.8\times 10^{-5}$ $2.7\times 10^{-5}$ $6.5\times 10^{-5}$ $5.9\times 10^{-5}$ $21.571072373$ $1.47\times 10^{-3}$
Table 3: Trends in the various quantities considered. The $tr_{x}$ and $tr_{y}$ trends, obtained from the JUP387 orbital theory, apply to both the orbit normal and spin axis coordinates $n/s_{x}$ and $n/s_{y}$, respectively. The $tr_{\alpha}$, $tr_{\delta}$ and $tr_{\mu}$ trends are obtained after the transformation described in Eqs. (11) and (24b). The rotation rate in $W$ is obtained from Eq. (33). The last column shows the small increment in $\delta_{S}$ due to the power reduction of a squared sine term in the second-order part of Eq. (54b). The number of digits retained in this table is related to the order of magnitude of the residuals from the quasi-periodic decomposition of JUP387, see Fig. 2.
Figure 3: Temporal evolution of the spin axis coordinates in the Laplace plane $s_{x}$ and $s_{y}$ for the four Galilean satellites for a solid satellite. The time coverage is different for each satellite. The purple point is the Laplace pole, for which by definition $s_{x}=s_{y}=0$. The black dots in the right graphs are the J2000 spin positions. We use here reconstructed series, which explains why we can extend our plot beyond the time interval covered by the orbital theory.
@@
In Table 4, we have considered a truncation threshold of the order of $0.0001^{\circ}$ on the $i_{j}$ terms, corresponding to the order of magnitude of the residuals in the orbit orientation angles between JUP387 and its frequency decomposition. We thus obtain $6$, $10$, 99, and $8$ terms for Io, Europa, Ganymede, and Callisto, respectively. The arguments (frequencies and phases) of the series result from the orbital dynamics and are combinations of the fundamental arguments associated with the Jovian system (see last columns of Table 4). However, arguments identified as being the same combination may have a slightly different value for each satellite, due to the fit to a numerical integration. The precision on the $i_{j}$, $\varepsilon_{j}$ and $\theta_{j}$ as given in Table 4 is $0.0001^{\circ}$ ($0.00001$ rad/y and $0.001^{\circ}$ in frequency and phase, respectively). The typical precision of the time-series for $s_{x/y}(t)$ reconstructed from Table 4 over the 190021001900-2100 interval is approximately $0.0003^{\circ}$ (see Table 5). This corresponds to a relative truncation error below $1\%$. Neglecting terms related to the secular trends $tr_{x/y}$ would lead to typical errors of about $0.003^{\circ}$ after $100$ years in $s_{x/y}(t)$, greater than the truncation error.
@@
Table 4: $(\Delta n_{x},\Delta n_{y})$ and $(\Delta s_{x},\Delta s_{y})$ series, according to the notations of Eqs. (49-51). Inclination, orbital and inertial obliquity amplitudes $i_{j}$, $\varepsilon_{j}$ and $\theta_{j}$ at the frequency and phase $f_{j}$ and $\varphi_{j}$ obtained from the JUP387 orbital theory and assuming that the satellites are entirely solid and rigid. These series can be used directly to reconstruct time series for $s_{x/y}(t)$, and indirectly to reconstruct the time series for the angles $\theta(t)$, $\varepsilon(t)$, $\psi(t)$, and $\zeta(t)$ (see Section 6.2). The phases $\varphi_{j}$ are given with respect to J2000 epoch. The last two columns show the identification to the fundamental arguments of Lainey et al. [2006] and the fundamental IAU WG frequencies $\dot{J}_{i}$. The number of digits retained in this table is related to the order of magnitude of the residuals from the quasi-periodic decomposition of JUP387, see Fig. 2.
@@
For each satellite, the orbital precession is dominated by the term with $j=1$ whose frequency corresponds to the node longitude proper frequency of that satellite. This may seem obvious, but note that this is not the case for all of the large satellites of Uranus, which strongly perturb each other [Baland et al., 2025]. The ratio $i_{2}/i_{1}$ is of $28\%$, $5\%$, $21\%$, and $12\%$, for Io, Europa, Ganymede, and Callisto by respectively. The dominant terms of the spin precession are here the same as the dominant terms of orbital precession, which will not necessarily be the case for satellites with an ocean, see Section 8.2. The amplitude $\varepsilon_{j}$ can reach a value such that $\theta_{j}$ is large or very small, depending on the ratio of $\omega_{f}/\dot{\Omega}_{j}$ with $\omega_{f}=\kappa/C$ the natural frequency of the free retrograde precession. The free precession mode periods are given in Table 2. In the limit case $\omega_{f}/\dot{\Omega}_{j}\rightarrow 0$ (relatively short forcing period), $\varepsilon_{j}$ tends towards $i_{j}$ and $\theta_{j}$ tends towards 0 (meaning that the spin axis tends to align with the Laplace pole, e.g. term $j=3$ for Callisto). For $\omega_{f}/\dot{\Omega}_{j}\rightarrow\pm\infty$ (relatively long forcing period), $\varepsilon_{j}$ tends towards 0 and $\theta_{j}$ tends towards $i_{j}$ (meaning that the spin axis tends to follow the precessing orbit normal, e.g. term $j=5$ for Io). When the ratio $\omega_{f}/\dot{\Omega}_{j}$ is close to 1-1, $\varepsilon_{j}$ and $\theta_{j}$ can be resonantly amplified (e.g. term $j=3$ for Ganymede). The ratio $\theta_{2}/\theta_{1}$ is of $26\%$, $5\%$, $18\%$, $16\%$ for Io, Europa, Ganymede and Callisto, respectively. Note that $\theta_{3}/\theta_{1}=22\%$ for Ganymede, while $i_{3}/i_{1}$ is only of $9\%$. The largest orbital obliquity amplitudes is $\varepsilon_{1}$ for Io, Europa and Callisto, but $\varepsilon_{3}$ for Ganymede. The second largest orbital obliquity amplitudes are $\varepsilon_{4}$, $\varepsilon_{5}$, $\varepsilon_{1}$, and $\varepsilon_{2}$ for Io, Europa, Ganymede, and Callisto, respectively.
@@
Table 5: Comparison of different errors affecting the angles over the 1900–2100 period. The first source of error arises from the truncation applied in Table 4. Each term in the $s_{x/y}$ series is truncated at a level of $0.0001^{\circ}$ in amplitude, which is also the truncation threshold applied to the series. The cumulative errors on $s_{x/y}(t)$, and therefore on $\theta(t)$, $\varepsilon(t)$, and $\zeta(t)$, lead to a total truncation error of about $0.0003^{\circ}$. The cumulative truncation errors on the node longitude $\psi$ reach up to $0.6^{\circ}$ for Io. The second source of error is the development error, arising when $\theta$, $\psi$, $\varepsilon$, and $\zeta$ are developed as series themselves. A third possible source of error stems from the omission of the linear (for $s_{x},s_{y}$, see Eq. 51) and Poisson terms (for $\theta$, $\psi$, $\varepsilon$, and $\zeta$, see Eqs. 53).